Tulsa Company has the following data available: 2X13 2X12 2X11 Sales (all credit sales) $800 $740 $675 Less Cost of Goods Sold 525 490 450 Gross Profit $275 $250 $225 What is the gross profit rate for Tulsa Company in 2X13? Has the gross profit rate improved or not improved since 2X12?

Select one: A. 52.4%, improved B. 52.4%, not improved C. 23.3%, improved D. 34.4%, not improved E. 34.4%, improved
Answer #1

Gross profit rate = Gross Profit /Sales

Gross profit in 2013 = 275/800 = 34.4%

Gross profit in 2012 = 250/740 = 33.8%

34.4% and increased

Option E